# Compact: Common-token Optimized Model Pruning Across Channels and Tokens

### v1

@@ 4.1 Experimental Setup
Baselines.
We compare with representative and state-of-the-art i) width pruning methods: SliceGPT, (Ashkboos et al., 2024), 2SSP (Sandri et al., 2025); ii) depth pruning methods: ShortGPT (Men et al., 2024), LaCo (Yang et al., 2024b).
LLMs to prune.
We evaluate on a diverse set of LLMs spanning architectures and scales: (i) Small LLMs: Qwen 2.5–0.5B, LLaMA 3.2–1B, and Gemma 3–1B. This mix covers three distinct families, enabling a robustness assessment across architectures; moreover, small LLMs are particularly challenging to prune, as they are often trained beyond the Chinchilla‐optimal compute–data balance, leaving limited redundancy. Nevertheless, pruning small LLMs is highly valuable for edge/on-device use and in privacy- or bandwidth-constrained settings (healthcare, classrooms, federated clients): it shrinks memory/storage, improves cache locality and end-to-end latency, lowers energy and thermal load, and reduces serving cost and carbon footprint. (ii) Large LLMs: LLaMA 3.1–8B and LLaMA 3.1–70B. Together with the 1B variant above, this suite evaluates pruning effectiveness across models differing by orders of magnitude in scale.
Evaluation tasks.
Following SliceGPT Ashkboos et al. (2024), we evaluate all pruning methods using 5 commonly-used natural language benchmarks: HellaSwag (HeSw) (Zellers et al., 2019), WinoGrande (WiGr) (Sakaguchi et al., 2020), ARC-C (Clark et al., 2018), ARC-E (Clark et al., 2018), and PIQA (Bisk et al., 2020). Since Jaiswal et al. (2024) shows that pruned LLMs have significantly degraded performance on more complex tasks, we add MMLU (Hendrycks et al., 2021) to test the models on more difficult general knowledge tasks, and GSM8K (Cobbe et al., 2021) to assess their performance on generation tasks. This collection enables us to obtain a more comprehensive overview of the model’s performance.
Evaluation Criteria.
Details about the evaluation setup and pruning hyperparameters can be found in Appendix A.2 and Appendix A.3, respectively.
  • We report the number of parameters (Para (M)) of each model after pruning, the percentage of parameters that were removed (Ratio (%)), the mean score of the 7 benchmarks (Avg), and the relative mean score compared to the unpruned dense model (Avg%). Note that SliceGPT does not report post-pruning parameter counts; we therefore report only the stated sparsity.
  • It is standard to evaluate pruned models on perplexity and downstream tasks. However, because we reduce vocabulary size, our perplexity naturally decreases, making comparisons to baselines unfair. Thus, we only report i) performance on downstream tasks, ii) efficiency regarding pruning time, inference time, and memory usage.
@@ 4.2.1 Results on smaller LLMs
The results are summarized in Table 2, which supports the following observations:
@@
The official implementations of our baseline pruning approaches support just a few model architectures. For instance, SliceGPT supported LLaMA, OPT, and Phi, LaCo supported LLaMA 2 and Baichuan, and ShortGPT only supported LLaMA. We needed to modify these implementations with architecture-specific changes in order to add support for more modern model families like LLaMA 3 and Qwen 2.5. This is because these architectures can differ significantly in how they handle self-attention, positional embeddings, layer normalization, and other details. The strength of Compact is that it only prunes the vocabulary embeddings and FFN blocks. These two model components have been standardized and remain unchanged across the vast majority of model architectures. As a consequence, Compact is architecture-agnostic and runs out of the box across diverse model families, requiring no model-specific modifications. This is most evident with the Gemma 3 family, which incorporates QK-norm and alternates local sliding-window and global attention layers—optimizations that prevented us from adapting baseline methods. Accordingly, baseline results for Gemma 3 are omitted in Table 2 (third block). In contrast, Compact operates on Gemma 3 without any changes.
Compact outperforms baselines with large margins.
Although prior work reports strong results on larger LLMs, the baseline methods perform poorly on smaller LLMs. On Qwen 2.5–0.5B, GSM8K accuracy collapses for all baselines at only a 10% pruning ratio. Likewise, at 10% pruning, MMLU drops to near-random for all models and baselines, with the sole exception of LaCo on Qwen 2.5–0.5B. Because MMLU and GSM8K are the most demanding tasks in our suite, these trends indicate that existing methods fail to preserve performance on challenging benchmarks for small LLMs. In contrast, Compact delays this collapse: it remains marginally above random on MMLU and GSM8K even at 35% pruning across all models. Across models and pruning ratios, Compact attains the highest mean score and leads on nearly all individual benchmarks, despite using a slightly higher pruning ratio than the baselines. Notably, at 35% pruning, Compact maintains similar performance to the baselines at 20%, demonstrating superior robustness under high compression.
Table 2: Pruning smaller LLMs (Qwen 2.5-0.5B, LLaMA 3.2–1B, and Gemma 3–1B) at a $\sim$10%, $\sim$20%, and $\sim$35% ratio. Please note pruning baselines cannot be applied to Gemma 3
@@
Method Para (M) Ratio (%) MMLU HeSw WiGr ARC-C ARC-E PIQA GSM8K Avg Avg% Qwen 2.5–0.5B Dense 494 0.00 47.3 52.2 56.4 32.3 58.2 69.9 34.9 50.2 100.0 Random - - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 57.0 ShortGPT 449 9.11 27.8 44.0 53.0 25.9 45.8 66.8 0.2 37.6 75.1 LaCo 449 9.11 46.1 45.5 56.3 28.2 51.6 65.5 0.4 41.9 83.6 SliceGPT - 10.71 23.2 43.2 53.3 26.4 50.3 63.8 0.0 37.2 74.1 2SSP 444 10.12 25.5 46.6 54.7 27.8 52.2 68.7 1.9 39.6 79.0 Compact 439 11.13 45.2 51.9 55.3 32.4 59.5 70.1 28.7 49.0 97.7 ShortGPT 405 18.02 25.0 37.7 52.0 27.1 41.7 62.1 0.0 35.1 70.0 LaCo 405 18.02 24.0 36.3 49.9 23.5 41.7 62.9 0.0 34.0 67.8 SliceGPT - 19.64 23.1 32.1 52.0 20.2 33.4 53.7 0.0 30.6 61.1 2SSP 397 19.64 24.3 40.9 53.8 25.3 43.9 64.3 0.5 36.1 72.0 Compact 394 20.24 44.1 48.1 55.4 30.6 53.3 66.6 26.3 46.3 92.4 ShortGPT 315 36.23 24.3 27.8 50.1 25.7 26.2 51.8 0.0 29.4 58.7 LaCo 315 36.23 23.9 28.2 47.9 23.9 30.6 56.2 0.0 30.1 60.0 SliceGPT - 36.61 23.1 29.0 51.5 22.5 30.7 53.4 0.0 30.0 59.9 2SSP 315 36.23 22.9 31.3 49.6 22.9 33.8 59.0 0.0 31.4 62.5 Compact 311 37.04 25.5 40.0 53.8 25.2 40.0 62.2 0.5 35.3 70.4 LLaMA 3.2–1B Dense 1236 0.00 36.6 63.8 60.7 36.2 60.7 74.5 5.4 48.3 100.0 Random - - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 59.2 ShortGPT 1112 10.03 23.4 48.0 60.2 30.4 49.8 68.0 0.0 40.0 82.8 LaCo 1112 10.03 24.4 48.4 52.0 29.5 47.9 69.3 0.5 38.9 80.5 SliceGPT - 10.16 23.1 49.4 54.0 28.6 43.4 64.0 0.0 37.5 77.7 2SSP 1114 9.87 30.5 55.5 57.9 32.9 56.7 72.9 3.0 44.2 91.6 Compact 1112 10.03 36.7 61.1 59.7 35.1 57.1 71.9 6.1 46.8 97.0 ShortGPT 993 19.66 22.8 40.0 55.3 29.9 35.2 58.9 0.0 34.6 71.7 LaCo 993 19.66 23.0 35.7 52.4 25.9 37.0 62.4 0.4 33.9 70.1 SliceGPT - 20.31 23.0 39.9 52.3 26.2 38.9 58.6 0.0 34.1 70.7 2SSP 993 19.66 26.8 46.9 53.9 27.1 50.4 68.1 2.2 39.3 81.5 Compact 989 19.98 30.6 54.4 58.6 32.0 51.3 69.9 3.1 42.8 88.8 ShortGPT 810 34.47 24.3 32.5 50.0 28.4 28.9 55.4 0.0 31.4 65.0 LaCo 810 34.47 23.2 37.3 50.1 23.9 29.3 55.3 0.0 29.9 61.9 SliceGPT - 35.16 23.0 30.4 51.2 22.0 32.9 53.4 0.0 30.4 63.1 2SSP 808 34.63 22.9 35.1 52.6 24.5 38.9 60.9 0.0 33.6 69.6 Compact 803 35.03 27.9 42.8 55.6 27.7 41.7 60.6 1.8 36.9 76.4 Gemma3–1B Dense 1009 0.00 24.9 62.1 59.0 38.2 71.9 74.8 2.4 47.6 100.0 Random - - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 60.0 Compact 908 10.01 24.9 60.3 59.0 39.1 69.0 74.1 1.7 46.9 98.4 807 20.02 25.0 55.4 59.0 37.9 63.3 70.0 1.7 44.6 93.6 656 34.99 24.2 45.1 55.9 26.5 46.4 65.3 0.5 37.7 79.2
@@ 4.2.2 Results on larger LLMs
LaCo failed to prune LLaMA 3.1-70B due to OOM errors, so it is omitted from our results. Our results are summarized in Table 3.
Compact is robust across scales via complementary pruning.
@@
We see that Compact achieves state-of-the-art performance for larger models as well, indicating that our method is highly robust to a wide range of model sizes. We attribute this robustness to our dual approach to width pruning. As model size increases, the proportion of vocabulary parameters decreases, which decreases the effectiveness of pruning vocabulary size. However, intermediate pruning becomes more effective as model size increases, similarly to most pruning methods (Xu et al., 2024). These two techniques complement each other’s strengths, leading to a robust hybrid pruning method. Although the performance gap between Compact and 2SSP narrows at the 70B size, we note that 2SSP’s hybrid approach of pruning FFN channels and entire attention blocks deviates from a transformer architecture, while Compact’s approach does not. This makes Compact more practical to deploy, while also achieving slightly higher performance.
Compact shows smooth, spike-free degradation.
@@
On LLaMA 3.1–8B, we observe an anomaly on MMLU: ShortGPT surpasses Compact at a 20% pruning ratio, but not at 10% or 35%, with its 20% score even exceeding its 10% score. This aligns with the step-like behavior reported by Gromov et al. (2025) for depth-pruning methods, wherein performance remains nearly intact up to a critical pruning threshold and then collapses abruptly. ShortGPT appears to exhibit this pattern, explaining the spike at 20%. In contrast, Compact (a width-pruning method) shows smooth MMLU degradation as pruning increases—hence the dip relative to ShortGPT at 20%, followed by recovery and a lead at 35%. A similar effect is seen on LLaMA 3.1–70B, where ShortGPT again outperforms on MMLU. Other benchmarks do not display step-like behavior, instead showing either smooth declines or immediate collapses. Notably, even at 35% pruning, Compact remains above random on both MMLU and GSM8K.
Table 3: Pruning larger LLMs (LLaMA 3.1–8B & LLaMA 3.1–70B) at a $\sim$10%, $\sim$20%, and $\sim$35% ratio.
Method Para (M) Ratio (%) MMLU HeSw WiGr ARC-C ARC-E PIQA GSM8K Avg Avg% LLaMA 3.1–8B Dense 8030 0.00 63.4 78.9 73.6 53.4 80.9 81.1 51.6 69.0 100.0 Random - - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 41.4 ShortGPT 7158 10.86 58.0 73.8 70.2 47.4 71.2 77.5 29.3 61.1 88.5 LaCo 7158 10.86 58.8 73.3 72.3 48.9 73.7 76.3 32.0 62.2 90.1 SliceGPT - 10.16 43.9 65.6 67.2 39.4 66.2 71.0 0.0 50.5 73.2 2SSP 7158 10.86 54.1 74.6 71.6 46.8 71.6 79.7 14.1 58.9 85.4 Compact 7227 10.00 59.6 75.2 73.7 50.3 74.9 78.4 27.8 62.9 91.1 ShortGPT 6503 19.02 58.6 64.9 68.4 42.2 58.3 71.6 0.6 52.1 75.5 LaCo 6503 19.02 24.1 54.0 55.3 29.2 51.1 72.4 0.4 40.9 59.3 SliceGPT - 20.12 24.5 51.4 61.8 30.3 49.0 61.9 0.0 39.8 57.8 2SSP 6425 19.99 37.4 67.2 68.4 38.1 61.8 76.8 4.3 50.6 73.3 Compact 6424 20.00 50.7 69.9 70.1 42.8 66.0 75.9 10.8 55.2 80.0 ShortGPT 5195 35.31 23.2 34.3 59.1 29.7 36.9 57.2 0.0 34.4 49.8 LaCo 5195 35.31 23.1 34.8 53.1 27.3 31.5 58.8 0.0 32.7 47.3 SliceGPT - 35.16 23.0 35.0 54.3 23.5 37.2 55.0 0.0 32.6 47.2 2SSP 5238 34.77 25.3 49.9 59.3 27.1 44.3 68.7 2.3 39.5 57.3 Compact 5220 34.99 35.9 56.0 63.3 30.8 48.4 70.6 1.7 43.8 63.5 LLaMA 3.1–70B Dense 70.6 0.00 75.2 85.0 79.5 64.7 86.7 84.4 80.6 79.4 100.0 Random - - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 36.0 ShortGPT 63.7 9.77 75.0 82.9 78.5 60.8 84.8 83.4 74.5 77.1 97.1 SliceGPT - 10.06 70.6 75.4 76.6 58.3 82.0 79.7 0.0 63.2 79.6 2SSP 63.6 9.92 73.4 84.7 78.6 63.9 85.4 84.2 75.1 77.9 98.1 Compact 63.5 10.06 73.5 84.5 79.5 64.0 86.0 84.1 76.0 78.2 98.5 ShortGPT 56.0 20.68 74.9 79.3 78.0 56.0 80.1 80.1 53.0 71.6 90.2 SliceGPT - 20.02 63.1 64.8 74.0 52.4 76.7 73.7 0.0 57.8 72.8 2SSP 56.4 20.11 68.8 83.7 76.2 59.5 82.1 83.7 62.1 73.7 92.8 Compact 56.4 20.11 70.6 83.3 76.2 59.7 82.6 83.7 62.6 74.1 93.3 ShortGPT 44.9 36.40 71.2 66.5 75.7 46.8 69.6 72.3 0.0 57.4 72.3 SliceGPT - 35.06 29.3 37.1 66.6 34.5 59.3 62.6 0.0 41.3 52.0 2SSP 45.8 35.13 58.5 77.7 72.1 50.7 74.2 81.6 25.0 62.8 79.1 Compact 45.9 34.99 59.6 78.1 72.8 53.2 76.1 81.3 25.0 63.7 80.2

### v2

@@ 4.1 Experimental Setup
Baselines. We compare with representative and state-of-the-art i) width pruning methods: SliceGPT, (Ashkboos et al., 2024), 2SSP (Sandri et al., 2025); ii) depth pruning methods: ShortGPT (Men et al., 2024), LaCo (Yang et al., 2024c). All methods use the default calibration set in their paper. For Compact, we use 256 calibration samples from the C4 dataset (Raffel et al., 2020).
LLMs to prune. We evaluate on a diverse set of LLMs spanning architectures and scales: (i) SLMs: Qwen 2.5–0.5B, LLaMA 3.2–1B, and Gemma 3–1B. This mix covers three distinct families, enabling a robustness assessment across architectures; moreover, small LLMs are particularly challenging to prune, as they are often trained beyond the Chinchilla‐optimal compute–data balance, leaving limited redundancy. Nevertheless, pruning small LLMs is highly valuable for edge/on-device use and in privacy- or bandwidth-constrained settings (healthcare, classrooms, federated clients): it shrinks memory/storage, improves end-to-end latency, lowers energy use, and reduces serving cost. (ii) LLMs: LLaMA 3.1–8B and LLaMA 3.1–70B. Together with the 1B variant above, this suite evaluates pruning effectiveness across a wide scale.
Evaluation tasks. Following SliceGPT (Ashkboos et al., 2024), we evaluate pruned models using HellaSwag (HeSw), WinoGrande (WiGr), ARC-C, ARC-E, and PIQA. Since Jaiswal et al. (2024) shows that pruned LLMs degrade more on complex tasks, we add MMLU for general knowledge, and GSM8K for generation tasks. This gives a more complete view of model performance.
Evaluation Criteria. Details about the evaluation setup and pruning hyperparameters can be found in Appendix A.2 and Appendix A.3, respectively. (A) We report the percentage of parameters that were removed (Ratio (%)), the mean score of the 7 benchmarks (Avg), and the relative mean score compared to the dense model (Avg%); (B) It is standard to evaluate pruned models on perplexity and downstream tasks. However, because we reduce vocabulary size, our perplexity naturally decreases, making comparisons to baselines unfair. Thus, we only report i) performance on downstream tasks, ii) efficiency regarding pruning time, inference time, and memory usage.
@@ 4.2 Performance on downstream tasks
Table 2: Pruning SLMs (Qwen 2.5-0.5B, LLaMA 3.2–1B, and Gemma 3–1B) at a $\sim$10%, $\sim$20%, and $\sim$35% ratio. Please note pruning baselines cannot be applied to Gemma 3
Method Ratio (%) MMLU HeSw WiGr ARC-C ARC-E PIQA GSM8K Avg Avg% Qwen 2.5–0.5B Dense 0.00 47.3 52.2 56.4 32.3 58.2 69.9 34.9 50.2 100.0 Random - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 57.0 ShortGPT 9.11 27.8 44.0 53.0 25.9 45.8 66.8 0.2 37.6 75.1 LaCo 9.11 46.1 45.5 56.3 28.2 51.6 65.5 0.4 41.9 83.6 SliceGPT 10.71 23.2 43.2 53.3 26.4 50.3 63.8 0.0 37.2 74.1 2SSP 10.12 25.5 46.6 54.7 27.8 52.2 68.7 1.9 39.6 79.0 Compact 11.13 45.2 51.9 55.3 32.4 59.5 70.1 28.7 49.0 97.7 ShortGPT 18.02 25.0 37.7 52.0 27.1 41.7 62.1 0.0 35.1 70.0 LaCo 18.02 24.0 36.3 49.9 23.5 41.7 62.9 0.0 34.0 67.8 SliceGPT 19.64 23.1 32.1 52.0 20.2 33.4 53.7 0.0 30.6 61.1 2SSP 19.64 24.3 40.9 53.8 25.3 43.9 64.3 0.5 36.1 72.0 Compact 20.24 44.1 48.1 55.4 30.6 53.3 66.6 26.3 46.3 92.4 ShortGPT 36.23 24.3 27.8 50.1 25.7 26.2 51.8 0.0 29.4 58.7 LaCo 36.23 23.9 28.2 47.9 23.9 30.6 56.2 0.0 30.1 60.0 SliceGPT 36.61 23.1 29.0 51.5 22.5 30.7 53.4 0.0 30.0 59.9 2SSP 36.23 22.9 31.3 49.6 22.9 33.8 59.0 0.0 31.4 62.5 Compact 37.04 25.5 40.0 53.8 25.2 40.0 62.2 0.5 35.3 70.4 LLaMA 3.2–1B Dense 0.00 36.6 63.8 60.7 36.2 60.7 74.5 5.4 48.3 100.0 Random - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 59.2 ShortGPT 10.03 23.4 48.0 60.2 30.4 49.8 68.0 0.0 40.0 82.8 LaCo 10.03 24.4 48.4 52.0 29.5 47.9 69.3 0.5 38.9 80.5 SliceGPT 10.16 23.1 49.4 54.0 28.6 43.4 64.0 0.0 37.5 77.7 2SSP 9.87 30.5 55.5 57.9 32.9 56.7 72.9 3.0 44.2 91.6 Compact 10.03 36.7 61.1 59.7 35.1 57.1 71.9 6.1 46.8 97.0 ShortGPT 19.66 22.8 40.0 55.3 29.9 35.2 58.9 0.0 34.6 71.7 LaCo 19.66 23.0 35.7 52.4 25.9 37.0 62.4 0.4 33.9 70.1 SliceGPT 20.31 23.0 39.9 52.3 26.2 38.9 58.6 0.0 34.1 70.7 2SSP 19.66 26.8 46.9 53.9 27.1 50.4 68.1 2.2 39.3 81.5 Compact 19.98 30.6 54.4 58.6 32.0 51.3 69.9 3.1 42.8 88.8 ShortGPT 34.47 24.3 32.5 50.0 28.4 28.9 55.4 0.0 31.4 65.0 LaCo 34.47 23.2 37.3 50.1 23.9 29.3 55.3 0.0 29.9 61.9 SliceGPT 35.16 23.0 30.4 51.2 22.0 32.9 53.4 0.0 30.4 63.1 2SSP 34.63 22.9 35.1 52.6 24.5 38.9 60.9 0.0 33.6 69.6 Compact 35.03 27.9 42.8 55.6 27.7 41.7 60.6 1.8 36.9 76.4 Gemma3–1B Dense 0.00 24.9 62.1 59.0 38.2 71.9 74.8 2.4 47.6 100.0 Random - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 60.0 Compact 10.01 24.9 60.3 59.0 39.1 69.0 74.1 1.7 46.9 98.4 20.02 25.0 55.4 59.0 37.9 63.3 70.0 1.7 44.6 93.6 34.99 24.2 45.1 55.9 26.5 46.4 65.3 0.5 37.7 79.2
@@ 4.2.1 Results on Smaller LLMs
Compact outperforms baselines by large margins.
Our results are summarized in Table 2. Although prior works report strong results on LLMs, they perform poorly on SLMs. On Qwen 2.5–0.5B, GSM8K accuracy collapses for all baselines at only 10% pruning. Likewise, at 10% pruning, MMLU drops to near-random for all models and baselines, with the sole exception of LaCo on Qwen 2.5–0.5B. Because MMLU and GSM8K are the most demanding tasks in our suite, these trends indicate that existing methods fail to preserve performance on challenging benchmarks for SLMs. In contrast, Compact delays this collapse: it remains marginally above random on MMLU and GSM8K even at 35% pruning across all models. Across models and pruning ratios, Compact attains the highest mean score and leads on nearly all individual benchmarks, despite using a slightly higher pruning ratio than the baselines. Notably, at 35% pruning, Compact maintains similar performance to the baselines at 20%, demonstrating superior robustness under high compression.
@@
The official implementations of existing approaches support just a few model architectures. For instance, SliceGPT supports LLaMA/OPT/Phi, LaCo supports LLaMA 2/Baichuan, and ShortGPT only supports LLaMA. Adding support for modern model families like LLaMA 3 and Qwen 2.5 required adding architecture-specific changes, since these architectures can differ significantly in how they handle self-attention, layer normalization, etc. The strength of Compact is that it only prunes the vocabulary embeddings and FFN blocks, which have been standardized and remain unchanged across the vast majority of model architectures. As a consequence, Compact is architecture-agnostic and runs out-of-the-box across many model families. This is most evident with Gemma 3, which uses QK-norm and alternating local and global attention layers—optimizations that prevented us from adapting our baselines. Accordingly, baseline results for Gemma 3 are omitted in Table 2. In contrast, Compact operates on Gemma 3 without any architecture-specific changes.
Table 3: Pruning larger LLMs (LLaMA 3.1–8B & LLaMA 3.1–70B) at a $\sim$10%, $\sim$20%, and $\sim$35% ratio. LaCo failed to prune LLaMA 3.1-70B due to OOM errors, so it is omitted from our results.
@@
Method Ratio (%) MMLU HeSw WiGr ARC-C ARC-E PIQA GSM8K Avg Avg% LLaMA 3.1–8B Dense 0.00 63.4 78.9 73.6 53.4 80.9 81.1 51.6 69.0 100.0 Random - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 41.4 ShortGPT 10.86 58.0 73.8 70.2 47.4 71.2 77.5 29.3 61.1 88.5 LaCo 10.86 58.8 73.3 72.3 48.9 73.7 76.3 32.0 62.2 90.1 SliceGPT 10.16 43.9 65.6 67.2 39.4 66.2 71.0 19.4 53.2 77.2 2SSP 10.86 54.1 74.6 71.6 46.8 71.6 79.7 14.1 58.9 85.4 Compact 10.00 59.6 75.2 73.7 50.3 74.9 78.4 27.8 62.9 91.1 ShortGPT 19.02 58.6 64.9 68.4 42.2 58.3 71.6 0.6 52.1 75.5 LaCo 19.02 24.1 54.0 55.3 29.2 51.1 72.4 0.4 40.9 59.3 SliceGPT 20.12 24.5 51.4 61.8 30.3 49.0 61.9 0.0 39.8 57.8 2SSP 19.99 37.4 67.2 68.4 38.1 61.8 76.8 4.3 50.6 73.3 Compact 20.00 50.7 69.9 70.1 42.8 66.0 75.9 10.8 55.2 80.0 ShortGPT 35.31 23.2 34.3 59.1 29.7 36.9 57.2 0.0 34.4 49.8 LaCo 35.31 23.1 34.8 53.1 27.3 31.5 58.8 0.0 32.7 47.3 SliceGPT 35.16 23.0 35.0 54.3 23.5 37.2 55.0 0.0 32.6 47.2 2SSP 34.77 25.3 49.9 59.3 27.1 44.3 68.7 2.3 39.5 57.3 Compact 34.99 35.9 56.0 63.3 30.8 48.4 70.6 1.7 43.8 63.5 LLaMA 3.1–70B Dense 0.00 75.2 85.0 79.5 64.7 86.7 84.4 80.6 79.4 100.0 Random - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 36.0 ShortGPT 9.77 75.0 82.9 78.5 60.8 84.8 83.4 74.5 77.1 97.1 SliceGPT 10.06 70.6 75.4 76.6 58.3 82.0 79.7 62.4 72.1 90.8 2SSP 9.92 73.4 84.7 78.6 63.9 85.4 84.2 75.1 77.9 98.1 Compact 10.06 73.5 84.5 79.5 64.0 86.0 84.1 76.0 78.2 98.5 ShortGPT 20.68 74.9 79.3 78.0 56.0 80.1 80.1 53.0 71.6 90.2 SliceGPT 20.02 63.1 64.8 74.0 52.4 76.7 73.7 0.0 57.8 72.8 2SSP 20.11 68.8 83.7 76.2 59.5 82.1 83.7 62.1 73.7 92.8 Compact 20.11 70.6 83.3 76.2 59.7 82.6 83.7 62.6 74.1 93.3 ShortGPT 36.40 71.2 66.5 75.7 46.8 69.6 72.3 0.0 57.4 72.3 SliceGPT 35.06 29.3 37.1 66.6 34.5 59.3 62.6 0.0 41.3 52.0 2SSP 35.13 58.5 77.7 72.1 50.7 74.2 81.6 25.0 62.8 79.1 Compact 34.99 59.6 78.1 72.8 53.2 76.1 81.3 25.0 63.7 80.2
@@ 4.2.2 Results on larger LLMs
Compact is robust across scales.
@@
Our results are in Table 3. We see that Compact achieves state-of-the-art performance for larger models as well, with over 80% performance at a 35% ratio, indicating that our method is highly robust to a wide range of sizes. We attribute this robustness to our dual approach to width pruning. As model size increases, the proportion of vocabulary parameters decreases, which decreases the effectiveness of pruning vocabulary size. However, intermediate pruning becomes more effective as model size increases, similarly to most pruning methods (Xu et al., 2024). These two techniques complement each other’s strengths, leading to a robust hybrid pruning method. Although the performance gap between Compact and 2SSP narrows at the 70B size, we note that 2SSP’s hybrid approach of pruning FFN channels and entire attention blocks deviates from the standard transformer architecture, while Compact’s approach does not. This makes Compact more practical to deploy, while also achieving slightly higher performance.
Compact shows smooth degradation.
@@
On LLaMA 3.1–8B, we observe that ShortGPT surpasses Compact at a 20% pruning ratio, but not at 10% or 35%, with its 20% score even exceeding its 10% score. This aligns with the step-like behavior in Gromov et al. (2025) for depth-pruning methods, where performance remains intact up to a critical threshold and then collapses abruptly. ShortGPT also exhibits this pattern, explaining the spike at 20%. In contrast, Compact (a width-pruning method) shows smooth MMLU degradation as pruning increases—hence the dip relative to ShortGPT at 20%, followed by recovery and a lead at 35%. Even at 35% pruning, Compact remains above random on both MMLU and GSM8K. A similar effect is seen on LLaMA 3.1–70B.
Table 4: Proportion of words retokenized after 35% pruning of Qwen 2.5–0.5B, by dataset.
Dataset Rare%
MMLU 4.43%
HellaSwag 3.48%
WinoGrande 5.01%
ARC-C 3.82%
ARC-E 3.95%
PIQA 5.68%
GSM8K 3.60%
C4 4.56%
Analysis: Why our pruning hurts performance minimally?
Changing the vocabulary affects how text is tokenized: If a token is removed during pruning, the text associated with the token is now tokenized as multiple shorter, more common tokens. We analyze how often rare tokens occur. We use the questions in each of our benchmarks, as well as 10k random samples from the C4 dataset. Our results are in Table 4. Our pruned model reduces vocabulary size from 150k to 50k, a 67% reduction. Despite this, only $4\%$ of words are tokenized differently from the original, regardless of text source. These results provide insight to the effectiveness of Vocab-pruning: significant proportions of vocabulary only affects a small fraction of text, so removing these vocabulary has little impact on performance.